Circular No. 6033

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) N. Walton, R. Catchpole, P. Rudd and A. Fitzsimmons, using the

Royal Observatories’ Isaac Newton Telescope at Roque de los Muchachos Observatory, report seeing the Na D lines in emission at the time of impact L. The emission first appeared on July 19.937 UT, reached maximum brightness on 19.938 and had vanished by 19.944. Other faint, as yet unidentified emission lines were also seen at the same time.

For the Hubble Space Telescope (HST) team, K. S. Noll, Space Telescope Science Institute, reports the following result: “We have identified CS_{2} in spectra of impact site G obtained with the HST Faint Object Spectrograph. At least 13 separate bands from 187.3 to 206.4 nm are clearly observed. In the same spectra we have previously positively identified S_{2} and enhanced NH_{3}, and we have tentative evidence for H_{2}S (cf. IAUC 6031). CS_{2} is an expected product of shock and/or solar extreme-ultraviolet induced chemical reactions between H_{2}S and CH_{4}. We have also searched for bands of SO and SO_{2}, but no band from either molecule has been observed. The two spectra were obtained on July 19.447 and 19.462 UT, more than three hours after the impact of fragment G. The 0′′.9 aperture was positioned at the center of the impact site.”

The South African Astronomical Observatory reports via the SL9 message center that J. Shykula used the 1.9-m telescope for high-speed K-band photometry of the impact of fragment S = 5. Definite brightening began around July 21.640 UT, with peak brightness around 21.645 and a return to relatively constant signal levels by 21.651; peak brightness was nearly twice that seen for impact Q, with a peak K ∼ 1.19. Using the 0.75-m telescope and PtSi camera K. Sekiguchi obtained images of Jupiter at 30-s intervals during the impact. A low-level brightening began around July 21.636 and began to fade before a much more dramatic brightening began around 21.640. Peak brightness occurred around 21.645, and by 21.653 there was definitely no remaining obvious excess brightness from this impact. The bright spot at the impact site could not easily be separated from those for impacts R, G and D. The S impact plume was considerably brighter than that seen for fragment H.

Observations with the MAGIC camera on the 3.5-m telescope at Calar Alto were affected by an electronic problem until very shortly before the S impact plume was detected at 2.3 μm on July 21.645 UT. It became brighter than the nearby spots of previous impacts within ∼ 2 min. It started fading on July 21.648 and eventually reached the level of the other bright spots.

1994 July 21 Brian G. Marsden